OpenStudy (jennychan12):

A concrete highway is built of slabs of 14 m long (20°C). How wide should the expansion cracks be (at 20°C) between the slabs to prevent buckling if the range of temperature is from -30°C to 50°C?

OpenStudy (jennychan12):

I know the equation is \[\Delta L = \alpha L \Delta T\]

OpenStudy (jennychan12):

is the ΔT 80°C? or is it 30°C?

OpenStudy (anonymous):

\[\Delta T\]is the difference in the temperature from the initial temp. (20C here) to the final temp.

OpenStudy (anonymous):

\[\Delta T=\text{max temp change}\\ \Delta T_1=T_{\rm final}-T_{\rm initial}=50-(20)=30^\circ C\\ \Delta T_2=T_{\rm final}-T_{\rm initial}=-30-(20)=-50^\circ C \]

OpenStudy (anonymous):

cracks should withstand both the temperature changes.

OpenStudy (anonymous):

i think we should only worry about the expansion. so the change in length due to the increase from the initial temperature should be taken care of...